OpenStudy (anonymous):

Solve the equation using quadrant formula. 9x^2 − 3x − 2 = 0

OpenStudy (anonymous):

-b/a = x1+x2 and c/a = x1*x2 3/9 = x1+x2 -2/9 = x1*x2

OpenStudy (anonymous):

Presumably you mean the quadratic formula. For an equation of the form: $$ax^2 + bx + c = 0,$$ The quadratic formula is: $$x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}$$ Your equation is: $$9x^2 - 3x -2 = 0.$$ In other words, a=9, b=-3, c = -2. Plug and play!